# Joint Distributed Generation Maximization and Radial Distribution Network Reconfiguration

### v1

@@ Abstract
This paper studies an optimization problem for joint radial distribution system network reconfiguration and power dispatch for distributed generation (DG) maximization. We provide counterexamples to show that for DG maximization, standard techniques such as interior point method (as in Matpower), linear approximation and second order cone relaxation (e.g., by Jabr et. al. and by Farivar and Low) do not deliver the desired control. Instead, we propose a control decision model based on exact DistFlow equations bypassing relaxation and a solution approach based on spatial branch-and-bound algorithm. We justify our work with comparative studies and numerical demonstrations with benchmarks and a 533-bus real example, performing reconfiguration and power dispatch in a time scale relevant for control center applications.
I Introduction
@@
Distribution networks are generally built in a looped fashion, but operated with one point normally-open in each loop resulting in a radial network. The set of switches to leave open (hence the network configuration) offers a degree-of-freedom that is normally not exploited in the industry. In this paper, we study the joint optimization of network reconfiguration and power control of distributed generation (DG), and we call this the reconfiguration optimal power flow (ROPF) problem, generalizing optimal power flow (OPF). For ROPF, we optimize an appropriate objective subject to network configuration and other common OPF constraints. The ROPF problem can model many active distribution system management applications (e.g., loss minimization, load balancing, renewable energy hosting capacity maximization [1, 2, 3, 4, 5, 6, 7]). However, ROPF is not as widely practiced as OPF in the industry, partially caused by the lack of efficient and reliable computation approaches to solve the problem. Therefore, the computation investigation of ROPF is the focus of this paper.
@@
The discrete decision variables associated with switch status make ROPF more difficult to solve than OPF. Diverse solution methods have been explored. For instance, heuristics and rule-based approaches have been considered due to their ease for online computation (e.g., [1, 7, 8]). To more explicitly search for the optimal solution, the ROPF problem is formulated and solved as an optimization problem, akin to OPF’s solution approach. A variety of methods have been proposed to solve the ROPF optimization problem. Meta-heuristics algorithms (e.g., [9, 10, 6]) are commonly used because of their modeling flexibility and relative ease of use. However, lack of optimality guarantee and unpredictable runtime are major drawbacks of these methods. Alternatively, more direct optimization methods are explored. For instance, [4] used the interior point method, and more recent research activities focus on ROPF optimization models in more specific classes so that more efficient algorithms can be applied. For example, mixed integer linear programming (MILP) models for ROPF and their extensions are extensively studied (e.g., [11, 12, 13, 5, 14]). Since the AC power flow equations are nonlinear, the MILP models are typically derived using some linear approximations of power system physics (e.g., DC power flow equations, linearized DistFlow equations). To balance model fidelity and computation tractability, mixed integer conic programming (MICP) models are proposed (e.g., [2, 15, 16, 17]). The MICP models are typically relaxations of the true ROPF problem in the sense that the feasible set of the former includes that of the latter. The relaxations are desired because they admit efficient solution algorithms often based on quadratic programming and second order cone programming. Successful stories of this approach are numerous (e.g., [2, 16, 18, 19]). However, a central question regarding MICP relaxation models is whether they are tight in the sense that the optimal solutions of the MICP models satisfy the constraints of the original ROPF problem. If not, the dispatch control based on the MICP models may violate operational constraints such as voltage and current limits. For OPF, guarantee of relaxation tightness can be obtained for certain “standard” objectives such as total line (active power) loss minimization, fuel cost minimization and conservation voltage reduction [15]. However, for ROPF, relaxation tightness is established only by numerical demonstrations. In response, this paper points out that for certain cases the MICP relaxation models are indeed not tight, and the resulting control will lead to violations of grid constraints. In particular, for DG maximization (related to renewable energy hosting capacity maximization [19] and curtailment minimization [20]) the MICP relaxation models from [2, 15] are found to be not tight. We support our claim with a simple three-bus counterexample showing all computation details. In addition, we reach the same conclusion when experimenting with ROPF instances derived from benchmarks from Matpower [21] and the REDS repository [7].
To complement the relaxation approach, we develop a ROPF model for radial distribution systems based on the (nonlinear) DistFlow equations [1]. Unlike [2, 15], we do not rely on relaxation and our model exactly describes the ROPF problem at the expense of being in a more difficult problem class: mixed integer program with linear constraints and bilinear equality constraints. To solve our model, we propose to use the spatial branch-and-bound (SBB) algorithm (e.g., [22]) as implemented in Gurobi (version 9 or later). SBB returns optimality gap certificates together with (sub)optimal solutions (if any). In addition, our experiment indicates that SBB is able to obtain high-quality solutions in reasonable amount of time for realistic benchmarks, indicating suitability for control center applications. For example, for a 533-bus real example, ROPF can be performed within 10 minutes when allowing up to four switch status changes.
@@
The rest of this paper is organized as follows: Section II introduces the notation, defines the ROPF optimization model and describes the proposed solution approach. Section III justifies the proposed ROPF model and its solution approach by comparing them with potential alternatives. Section IV illustrates the practical value of the proposed work by considering the renewable energy hosting capacity maximization of a 533-bus distribution network in Sweden.
II Problem Description
II-A Problem Overview
@@
We consider balanced power distribution systems with line shunt elements ignored. The lines are equipped with switches which can be closed or open. The network is built with loop(s) but operates in a radial topology. Power exchange is possible at the slack bus. In addition, load and DG are present in the system. The loads are assumed to be known. The DG sources are interfaced with inverters where both active and reactive power outputs are considered adjustable. By jointly adjusting network configuration and inverter outputs, the ROPF problem considered in this paper seeks to maximize the total distributed active power generation in the system with the following constraints: (a) power system physics (DistFlow equations), (b) network radiality and allowable switch status changes, (c) inverter output bounds, apparent power capacity, power factor and (d) grid constraints (voltage and line current limits).
II-B Notation and Assumptions
@@
  • For any bus $n\in\bar{\mathcal{N}}$, the voltage is $v_{n}$ and we define $\nu_{n}:=|v_{n}|^{2}$ as the squared magnitude voltage for bus $n$. $\nu_{n}$ is a state decision variable in the ROPF problem.
@@
  • In a radial network, each bus $n\in\mathcal{N}$ is associated with exactly one line, denoted by line $n$, with reference direction pointing away from bus 0. The current and power flow on the line follow the same reference direction of the line.
@@
  • Line shunt elements are ignored. For any bus $n\in\mathcal{N}$, $\ell_{n}$ is the squared magnitude current from bus $\pi_{n}$ to bus $n$ (also the value from $n$ to $\pi_{n}$). $\ell_{n}$ is a state decision variable in the ROPF problem.
  • For a line $\{n,k\}$ in the reconfigurable network, the series resistance and reactance are $r_{nk}$ and $x_{nk}$ respectively. In addition, the square line current upper limit is denoted by ¯$\overline{\ell}_{nk}$. Note that $r_{nk}=r_{kn}$, $x_{nk}=x_{kn}$ and ¯$\overline{\ell}_{nk}=\overline{\ell}_{kn}$.
@@
II-C ROPF Problem Constraints Derivation
II-C1 DistFlow Equations for Fixed Network Configuration
@@
where $r_{n}$ and $x_{n}$ are the series resistance and series reactance of line $n$, respectively. The set $c(n)$ is defined in Section II-B.
II-C2 Network Configuration Modelling
We impose bounds on the switch status decision variables $\alpha_{nk}$ defined in Section II-B to indicate always closed and always open switches:
@@ II-C2 Network Configuration Modelling
A network with $N+1$ buses is radial if and only if it is connected and contains exactly $N$ lines. Radiality can be imposed by the (necessary and sufficient) single-commodity flow conditions. Using the parent decision variables $\pi_{nk}$ defined in Section II-B and auxiliary continuous “flow” variables $f_{ij}$ for $(i,j)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$, these conditions can be described as
@@
(3a) and (3b) specify that one unit of (fictitious) flow should be shipped from the slack bus to each non-slack bus through a directed network whose edges $(n,k)$ satisfy $\pi_{nk}=1$. This condition is the same as network connectedness. Also, (3c) specifies that the network has exactly $N$ edges. Hence, (3) is necessary and sufficient for network radiality. Other conditions guaranteeing network radiality include multi-commodity flow, Miller-Tucker-Zemlin (MTZ), Dantzig-Fulkerson-Johnson (DFJ), Martin [23]. See [24, Section IV-C] for the necessity to properly impose the radiality constraints.
@@
Note that when a switch is closed another switch must be opened simultaneously to keep the radial network configuration. Therefore, $K$ is typically chosen as an even number.
II-C3 DistFlow Equations for Configurable Network
When the network configuration is parameterized by decision variable $\pi$ (also $\alpha$), certain configuration related terms in (1) should be generalized. For instance, the children set $c(n)$ in (1a) and (1b) is not determined before the problem is solved. Thus, the following modifications are needed in (1a) and (1b)
@@ II-C3 DistFlow Equations for Configurable Network
to describe the sum of power flows leaving bus $n\in\mathcal{N}$. Similarly, while line $n$ always points to bus $n$, the parent $\pi_{n}$ (hence the actual line) is not fixed before optimization. The line parameters and parent bus voltage in (1a) to (1d) become
@@
where $r_{kn}$ and $x_{kn}$ are the series resistance and reactance of line $\{n,k\}$ as defined in Section II-B. Hence, the DistFlow equations for configurable networks become
@@
Contrary to the fixed configuration setting in (1), the line variables $P_{n}$, $Q_{n}$ and $\ell_{n}$ in (5) do not correspond to a predefined line in the network. These line variables can be associated with any one of the lines incident to bus $n$. The actual association is determined by the values of the $\pi$ variables after the optimization run.
In (5a), the bilinear term $\pi_{nk}P_{k}$ with 0-1 binary variable $\pi_{nk}$ and continuous variable $P_{k}$ can be “linearized” without incurring any approximation. Suppose the continuous variable $P_{k}$ is bounded such that $\underline{P}_{k}\leq P_{k}\leq\overline{P}_{k}$. By introducing auxiliary continuous variable $z$ satisfying
@@
the bilinear term $\pi_{nk}P_{k}$ can be replaced by $z$. The claim can be verified as follows: when $\pi_{nk}=0$ the bilinear term $\pi_{nk}P_{k}$ should be zero. In this case, (6a) forces $z$ to be zero while (6b) holds trivially as $P_{k}-\overline{P}_{k}\leq 0\leq P_{k}-\underline{P}_{k}$. On the other hand, when $\pi_{nk}=1$ the bilinear term $\pi_{nk}P_{k}$ should be equal to $P_{k}$. In this case, (6a) holds trivially while (6b) specifies that $z=P_{k}$. Other binary-continuous bilinear terms $\pi_{kn}\ell_{n}$, $\pi_{nk}Q_{k}$ and $\pi_{kn}\nu_{k}$ in (5) can be linearized similarly. However, in (5d) the terms $(\pi_{kn}\nu_{k})\ell_{n}$ remain bilinear even after the linearization of $\pi_{kn}\nu_{k}$. This complicates the solution process of the ROPF problem.
II-C4 Inverter Operational Constraints
The inverter operational constraints are divided into three groups: (a) generation bounds, (b) apparent power rating, and (c) minimum power factor. These requirements are
@@ II-C5 Grid Constraints
where ¯$\overline{\ell}_{kn}$ is the given square current upper limit for line $\{k,n\}$. Typically, ¯$\sqrt{\overline{\ell}_{kn}}$ is up to a few hundred amperes.
The linearization of the bilinear terms $\pi_{nk}P_{k}$ in (5a) and $\pi_{nk}Q_{k}$ in (5b) requires variable bounds on the line power flows $P_{k}$ and $Q_{k}$. If these bounds are not provided, they can be derived as follows:
@@
where $\nu^{U}_{k}:=\max\limits_{(n,k)\mid\overline{\alpha}_{nk}=1}\overline{\nu}_{n}$ and ${\ell}^{U}_{k}:=\max\limits_{(n,k)\mid\overline{\alpha}_{nk}=1}\overline{\ell}_{nk}$. The bounds in (10) and (11) follow necessarily from (8) and (9). $P_{k}+jQ_{k}=v_{\pi_{k}}(I_{k})^{*}$ with $v$ and $I$ being the complex voltage and line current. Thus, $|P_{k}|\leq|P_{k}+jQ_{k}|\leq|v_{\pi_{k}}||I_{k}|\leq\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}}$. This establishes (10). The same argument shows (11).
II-D DG maximizing ROPF Model and Solution Approach

### v2

@@ Abstract
Network reconfiguration can significantly increase the hosting capacity (HC) for distributed generation (DG) in radially operated systems, thereby reducing the need for costly infrastructure upgrades. However, when the objective is DG maximization, jointly optimizing topology and power dispatch remains computationally challenging. Existing approaches often rely on relaxations or approximations, yet we provide counterexamples showing that interior point methods, linearized DistFlow and second-order cone relaxations all yield erroneous results. To overcome this, we propose a solution framework based on the exact DistFlow equations, formulated as a bilinear program and solved using spatial branch-and-bound (SBB). Numerical studies on standard benchmarks and a 533-bus real-world system demonstrate that our proposed method reliably performs reconfiguration and dispatch within time frames compatible with real-time operation.
1 Introduction
@@
Distribution networks are generally built in a looped fashion, but operated with one point normally-open in each loop resulting in radial networks. The set of switches to leave open (hence the network configuration) offers a degree-of-freedom that is normally not exploited in the industry. In this paper, we study the joint optimization of network reconfiguration and power control of distributed generation (DG), and we call this the reconfiguration optimal power flow (ROPF) problem, generalizing optimal power flow (OPF). While enforcing network configuration and operational constraints (e.g., voltage and line current limits and DG restrictions), the ROPF problem traditionally seeks to minimize objectives including total line loss, fuel cost or voltage deviation (e.g., [3, 15, 33, 1, 24, 2]). However, in this paper we consider the maximization of total DG output in ROPF as an emerging topic for renewable energy integration (e.g., [5, 10, 27, 13]). The DG maximization objective has a crucial effect on the tractability of the ROPF problem and the usefulness of the resulting dispatch solution. This is the first main discovery of this paper.
@@
Diverse solution methods for the ROPF problem have been explored. For instance, heuristics and rule-based approaches have been considered due to their ease for online computation (e.g., [3, 1, 12]). To more explicitly search for the optimal solution, the ROPF problem is formulated and solved as an optimization problem, akin to OPF’s solution approach. A variety of methods have been proposed to solve the ROPF optimization problem. Meta-heuristics algorithms (e.g., [32, 35, 10]) are commonly considered due to their flexibility and ease of use. However, lack of optimality guarantee and unpredictable runtime are major drawbacks of these methods. Thus, more direct optimization methods are explored. For example, using DC power flow equations or linearized DistFlow equations to approximate the power system physics, [23, 9, 6, 26] investigate mixed integer linear programming (MILP) models for ROPF. Alternatively, to directly include the AC power flow equations, mixed integer conic programming (MICP) ROPF models are proposed (e.g., [15, 8, 33, 17]). Theoretical support and numerical demonstrations have been reported (e.g., [8, 20, 15, 33, 19, 13]). However, it is not well recognized that the objective of the ROPF problem plays an important role in the suceess of the solution approaches. In particular, except for [13] all aformentioned previous works consider “classical” objectives such as loss minimization, convex quadratic fuel cost minimization, conservation voltage reduction, etc. For the “non-classical” objective of DG maximization in this paper, it turns out that popular methods such as the MICP apporach and interior point algorithm may lead to undesirable control dispatch violating grid constraints such as voltage and current limits. We support our claim with a repeatable case study involving a three-bus example with all computation details presented. To the best of our knowledge, the adverse effects of these pitfalls have not been well recognized in the literature.
The difficulty with DG maximizing ROPF necessitates a solution method with acceptable accuracy and efficiency. With the only nonlinearity in the model being polynomials (due to the power flow equations), the ROPF problem can be interpreted as a bilinear programming problem using specialized solver (e.g., Gurobi). Similar approaches have been studied in [18, 5], where the ROPF problem is treated as a more general mixed integer nonlinear programming (MINLP) problem and solved using the spatial branch-and-bound (SBB) algorithm. In terms of accuracy, the SBB algorithm is sufficient since it guarantees optimality regardless of the ROPF objective type. However, since SBB is fundamentally an exhaustive search it is important to investigate its computational efficiency for the DG maximizing ROPF problem. The rather long runtime reported in [18, 5] does not imply efficiency especially for real-time applications. The efficiency of SBB is significantly affected by the choice of problem formulation and algorithm implementation. A crucial factor in problem formulation is the model of power system physics described by the power flow equations. Typical forms of the power flow equations include, for example, AC power flow equations with voltages in polar coordinates (e.g., [18]) and voltages in rectangular coordinates (e.g., [5]), branch flow equations [8], DistFlow equations [3], modified AC power flow equations (e.g., [7, 16]), etc. On the other hand, different SBB implementations (e.g., Gurobi, BARON, Knitro) behave differently because they make different assumptions about the problem structure. Through our case studies, we identify the combination of DistFlow equations and Gurobi (version 9 or later, with recently added capability to handle bilinear programs) as a promising choice of solution method for the DG maximizing ROPF problem. For instance, for a 533-bus real example in Sweden, ROPF can be performed within 10 minutes when allowing up to four switch status changes. To the best of our knowledge, successful demonstration of SBB to solve ROPF models with the exact DistFlow equations is not well-known, even though the DistFlow equations are commonly relaxed and used in the MICP approach (e.g., [8]). We emphasize that the appropriate choice of formulation and SBB solver is important. We will provide counterexamples to illustrate the poor computational performance of some improper choices. We will also demonstrate that the DistFlow equations play a vital role in the efficient computation of the ROPF problem. This is the second main discovery of this paper.
The main novelties and contributions of this paper are listed as follows:
  • We raise the awareness that the DG maximizing objective fundamentally changes the tractability of the ROPF problem and illustrate the pitfalls of possible overextension of popular solution methods such as MICP relaxation, interior point algorithms and linearization approximation.
  • We point out the importance of the choice of problem formulation and optimization solver for the DG maximizing ROPF problem, by illustrating a promising choice and other improper choices.
@@
The rest of this paper is organized as follows: Section 2 introduces the notation, defines the ROPF optimization model and describes the proposed solution approach. Section 3 illustrates and evaluates the performance of the proposed ROPF approach. Section 4 justifies the proposed method by showing pitfalls encountered by common alternatives for DG maximizing ROPF. Section 5 discusses a practical renewable energy hosting capacity maximization problem for a 533-bus distribution network in Sweden using the proposed method.
2 Problem Description
2.1 Problem Overview
@@
We consider balanced power distribution systems with line shunt elements ignored so that the system can be described by the DistFlow equations [3]. The lines are equipped with switches which can be closed or open. The network is built with loop(s) but is operated in a radial topology. Power exchange is possible at the slack bus. In addition, load and DG are present in the system. The loads are assumed to be known. The DG sources are interfaced with inverters where both active and reactive power outputs are considered adjustable. By jointly adjusting network configuration and inverter outputs, the ROPF problem considered in this paper seeks to maximize the total distributed active power generation in the system with the following constraints: (a) DistFlow equations, (b) network radiality and allowable switch status changes, (c) inverter power output bounds, apparent power capacity, power factor and (d) grid constraints (voltage and line current limits).
We comment that our adopted DistFlow model, while restricted by the three-phase balance assumption, is well accepted as evidenced by numerous influential results built upon the same assumption (e.g., [3, 15, 33]). Also, it is common to find representative distribution system models described exactly by the DistFlow equations (e.g., MATPOWER [36], REDS repository [1], EU examples [22]).
2.2 Notation and Assumptions
@@
  • For any bus $n\in\bar{\mathcal{N}}$, the voltage is $v_{n}$. We define $\nu_{n}:=|v_{n}|^{2}$ as the squared voltage for bus $n$. $\nu_{n}$ is a state decision variable in the ROPF problem.
@@
  • In a radial network, each bus $n\in\mathcal{N}$ is associated with exactly one line, denoted by line $n$, with reference direction pointing away from bus 0. The current and power flow follow the same reference direction of the line.
@@
  • Line shunt elements are ignored. For any bus $\in\mathcal{N}, n\ell_{n}$ is the squared magnitude current from bus $\pi_{n}$ to bus $n$ (also the value from $n$ to $\pi_{n}$). $\ell_{n}$ is a state decision variable in the ROPF problem.
  • For a line $\{n,k\}$ in a reconfigurable network, the series resistance and reactance are $r_{nk}$ and $x_{nk}$ respectively. Also, the square line current upper limit is denoted by $\overline{\ell}_{nk}$. Note that $r_{nk}=r_{kn}$, $x_{nk}=x_{kn}$ and $\overline{\ell}_{nk}=\overline{\ell}_{kn}$.
@@
2.3 ROPF Problem Constraints Derivation
2.3.1 DistFlow Equations for Fixed Network Configuration
@@
where $r_{n}$ and $x_{n}$ are the series resistance and series reactance of line $n$, respectively. The set $c(n)$ is defined in Section $2.$
2.3.2 Network Configuration Modelling
We impose bounds on the switch status decision variables $\alpha_{nk}$ defined in Section 2.2 to indicate always closed and always open switches:
@@ 2.3.2 Network Configuration Modelling
A network with $N+1$ buses is radial if and only if it is connected and contains exactly NN lines. Radiality can be imposed by the (necessary and sufficient) single-commodity flow conditions. Using the parent decision variables $\pi_{nk}$ defined in Section 2.2 and auxiliary continuous “flow” variables $f_{ij}$ for $(i,j)\in\bar{\mathcal{N}}\times\bar{\mathcal{N}}$, these conditions can be described as [28, 29]
@@
Note that when a switch is closed another switch must be opened simultaneously to keep the radial network configuration. Hence, $K$ is typically an even number.
2.3.3 DistFlow Equations for Configurable Network
When the network configuration is parameterized by decision variable $\pi$ (also $\alpha$), certain configuration related terms in (1) should be generalized. For instance, the children set $c(n)$ in (1a) and (1b) is not determined before the problem is solved. Thus, the following changes are needed in (1a) and (1b)
@@ 2.3.3 DistFlow Equations for Configurable Network
to describe the sum of power flows leaving bus $n\in\mathcal{N}$. Similarly, while line $n$ always points to bus $n$, the parent $\pi_{n}$ is not fixed before optimization. The line parameters and parent bus voltage in (1a) to (1d) become
@@
where $r_{kn}$ and $x_{kn}$ are the series resistance and reactance of line $\{n,k\}$ as defined in Section $2.$ Hence, the DistFlow equations for configurable networks become
@@
Contrary to the fixed configuration setting in (1), the line variables $P_{n}$, $Q_{n}$ and $\ell_{n}$ in (5) do not correspond to a predefined line in the network. These line variables can be associated with any line incident to bus $n$. The actual association is determined by the values of the $\pi$ variables at optimality. We note that the formulation in (5) is different from standard reconfiguration-type formulation in the literature, where big-M type status changes are directly imposed on (5c).
In (5a), the bilinear term $\pi_{nk}P_{k}$ with 0-1 binary variable $\pi_{nk}$ and continuous variable $P_{k}$ can be equivalently replaced by an auxiliary variable with additional linear constraints as follows. Suppose the continuous variable $P_{k}$ is bounded such that $\underline{P}_{k}\leq P_{k}\leq\overline{P}_{k}$. With auxiliary continuous variable $z$ satisfying
@@
the bilinear term $\pi_{nk}P_{k}$ can be replaced by $z$. The claim can be verified by separately considering the cases when $\pi_{nk}=0$ and $\pi_{nk}=1$. Other binary-continuous bilinear terms $\pi_{kn}\ell_{n}$, $\pi_{nk}Q_{k}$ and $\pi_{kn}\nu_{k}$ in (5) can be replaced by auxiliary variables in a similar fashion. However, in (5d) the terms $(\pi_{kn}\nu_{k})\ell_{n}$ remain bilinear even after the replacement of $\pi_{kn}\nu_{k}$. This complicates the solution process of the ROPF problem.
2.3.4 Inverter Operational Constraints
The inverter operational constraints include (a) generation bounds, (b) apparent power rating, and (c) minimum power factor. These constraints are
@@ 2.3.5 Grid Constraints
where $\overline{\ell}_{kn}$ is the given square current upper limit for line $\{k,n\}$. Typically, $\sqrt{\overline{\ell}_{kn}}$ is up to a few hundred amperes.
The replacement of the bilinear terms $\pi_{nk}P_{k}$ in (5a) and $\pi_{nk}Q_{k}$ in (5b) requires variable bounds on the line power flows $P_{k}$ and $Q_{k}$. If these bounds are not provided, they can be derived as follows:
@@
where $\nu^{U}_{k}:=\max\limits_{(n,k)\mid\overline{\alpha}_{nk}=1}\overline{\nu}_{n}$ and ${\ell}^{U}_{k}:=\max\limits_{(n,k)\mid\overline{\alpha}_{nk}=1}\overline{\ell}_{nk}$. The bounds in (10) and (11) follow necessarily from (8) and (9). $P_{k}+jQ_{k}=v_{\pi_{k}}(I_{k})^{*}$ with $v$ and II being the complex voltage and line current. Thus, $|P_{k}|\leq|P_{k}+jQ_{k}|\leq|v_{\pi_{k}}||I_{k}|\leq\sqrt{\nu^{U}_{k}{\ell}^{U}_{k}}$. This establishes (10). The same argument shows (11).
2.4 DG maximizing ROPF Model and the Proposed Solution Approach